# Multi-Modal Multi-Task (M3T) Federated Foundation Models for Embodied AI: Potentials and Challenges for Edge Integration

### v1

@@ III Use Cases of FFMs in Embodied AI and EMBODY Dimensions
Use Cases of FFMs in Embodied AI: We next provide three examples on the use cases of FFMs in embodied AI. (1) FFMs can empower smart factories, where cooperative robots with diverse sensors, actuators, and policies adapt to dynamic workflows and human collaborators. Unlike traditional FMs that may need centralized retraining, FFMs enable decentralized training/adaptation, preserving proprietary data and meeting strict privacy regulations. (2) In domestic environments, FFMs enable assistive robots to learn and personalize to users’ routines, preferences, and spaces without exposing sensitive household data, overcoming a key limitation of FM-only solutions. (3) In immersive XR systems, FFMs enable distributed headsets and wearables to collaboratively fine-tune models for body-language, gaze, and gesture recognition, achieving real-time on-device personalization.
Unlike conventional FMs, which are frozen after centralized pretraining and often lack situational responsiveness, FFMs unlock continuous, privacy-preserving adaptation to new users, tasks, and hardware configurations in all the above scenarios.
Articulation of EMBODY Dimensions: In the above scenarios, the EMBODY dimensions naturally emerge, underscoring the critical challenges that naive implementations of FFMs in embodied AI cannot address.
@@ IV Open Research Directions: An EMBODY-Aligned Agenda for FFMs
We next revisit the aforementioned EMBODY dimensions, aiming to tailor a series of research directions to address them. Each part introduces a core challenge and proposes open research directions grounded in the modular FFM framework. The directions are intentionally framed at a high-level to allow for diverse interpretations and encourage innovation in this underexplored domain.
Figure 3: Visualizations of the envisioned future research directions. (a): On-Device Continual Learning. Left Plot: Drift-Aware Module Aggregation. Right Plot: Temperature-Aware Model Aggregations. (b): Distributed Control and Autonomy. Left Plot: Federated Multi-Agent Contrastive Alignment. Right Plot: Decentralized MoE Federation. (c): Yielding Safety, Privacy, and Personalization. Left Plot: Module-Aware Differential Privacy. Right Plot: Federated Behavioral Shadowing Buffers.
@@ IV-A Embodiment Heterogeneity
The embodied heterogeneity, caused by agents operating in various physical environments, poses a challenge to building shared FFMs that generalize across embodiments (e.g., can be trained on one or more types of robots and still perform effectively when deployed on others). To address this, we envision the following (see Fig. 2(a)):
(i) Embodiment-Conditioned Expert Routing: Modality-specific MoME layers can be augmented with learned embodiment tokens that encode hardware traits (e.g., kinematic range, actuator count, sensor precision). These tokens can then act as soft keys for expert selection, enabling information routing decisions inside the model that reflect physical context during inference and model personalization.
(ii) Cross-Embodiment Consistency Regularization: During federated training, agents with different embodiments often diverge in representation space (i.e., variations in hardware, sensing, or control may lead to different internal encodings of similar tasks). To address this, we envision embedding a consistency loss between related tasks performed by heterogeneous embodiments to make a closer connection between similar tasks performed across embodiment shifts (i.e., changes in the physical configuration of agents, such as differences in arm length, gripper type, or sensor placement, which affect how the same task is perceived and executed).
@@ IV-B Modality Richness and Imbalance
To handle modality richness and imbalance of agents, caused by the (temporal) variations of their sensory inputs, we envision the following (see Fig. 2(b)):
(i) Modality-Aware Clustered Training: We envision agent clustering techniques that group agents with similar tasks while ensuring the combined modalities in each group/cluster cover broader modalities than any single agent. Within each cluster, the server can perform federated aggregation across selected MoME experts, enabling agents to learn not only from peers with similar tasks but also from complementary modalities they lack. Crucially, in each cluster, experts within the MoME layer of one agent can also learn the hidden structures and relationships between modalities by leveraging knowledge from agents with richer modality sets.
(ii) Cross-Modal Recovery via Expert Cascades: To improve robustness against missing modalities, we envision expert cascades within the MoME layer: a series of specialized models that activate only when certain types of data are missing or low-quality. These experts should be trained in a federated manner, where agents locally train their experts using their available sensor data and send only the expert model updates to the server. The server can then aggregate these updates for each of these experts across agents, allowing knowledge sharing without exposing raw data. This way, agents can collaboratively learn to fill in missing modalities (e.g., predicting touch input from vision and movement sensors) while keeping the training decentralized and privacy-preserving.
@@ IV-C Bandwidth and Compute Constraints
To consider the fact that many embodied agents operate on edge processors with strict energy/compute and bandwidth limitations, we envision the following (see Fig. 2(c)) :
(i) Context-Aware Federated Caching: We envision a context-aware edge caching mechanism for FFMs, where the edge server caches the modules received from the agents during the model aggregations, adhering to its memory constraints. Here, agents participating in FFM training first query their needed modules from their nearest server based on their current task and environment (e.g., “following the user in hallway”) to avoid redundant local model updates and transmissions. If another agent has contributed a similar module to the edge cache during previous federated rounds, the querying agent can download and reuse that module with only a light fine-tuning. This mechanism can leverage semantic task signatures (e.g., compact embeddings from recent observation-action histories) to match cached modules with the agent’s current context.
(ii) Event-Triggered Scheduling of FFM Modules: In FFMs, we envision that event-triggered scheduling can reduce unnecessary computation/communication. Specifically, rather than relying on fixed periodic model transmissions to the server, agents can dynamically trigger the training of specific FFM modules (e.g., encoders, adapters, or expert layers) only when significant context/environment changes, task shifts, or performance drops are detected. For example, a mobile robot might update its navigation task head only when entering a new environment or encountering unfamiliar obstacles.
@@ IV-D On-Device Continual Learning
To handle on-device continual learning of embodied agents, caused by facing evolving environments, user preferences, and task definitions — see Fig. 3(a) — we envision the following:
(i) Drift-Aware Module Creation: We envision task drift management for FFMs. At the agent-level, task drift may indicate the need to fine-tune or split a local task head into two. For example, in the case of “cleaning” task for a home robot, given the variations in room layouts, it may benefit from separate heads for “kitchen cleaning” and “bedroom cleaning.” The key challenge lies in determining whether this drift is isolated/local or systematic/global. If drift patterns are detected across multiple agents (e.g., via cluster-level model divergence metrics), the server may create a new global task head. Otherwise, agent-level fine-tuning (e.g., cross-task proximal terms or adding adapters) remains the preferred path to avoid global model bloat and maintain its generalization.
(ii) Temperature-Aware Model Aggregations: In dynamic FL environments, where data and modality drift are common, different model modules can be at varying levels of readiness or maturity — some are well-trained and performing reliably, called “warm”; while some are newly introduced or still adapting (e.g., a freshly added encoder for an emerging modality), called “cold”. To account for these variations, we envision a temperature-aware aggregation strategy, where the FL server adjusts the weight or frequency of modules’ aggregation based on their current “temperature.” Well-performing, stable modules can be sent for aggregation less frequently to avoid marginal performance gains, while newer or less confident modules receive more frequent and aggressive aggregation to accelerate their performance across agents.
@@ IV-E Distributed Control and Autonomy
To enable embodied agents to adapt to their local tasks and contexts in a coordinated yet decentralized manner — see Fig. 3(b) — we envision the following:
(i) Federated Multi-Agent Contrastive Alignment: Embodied agents often can perform similar tasks while operating in different environments, using different hardware, or facing different user preferences. We envision a shared contrastive learning objective that works through teaching agents to pull the internal representations (latent spaces) of similar tasks closer and push apart those of non-similar tasks. In our case, this will align the task understanding of agents with similar tasks without sharing raw data. For example, an aerial drone and a ground robot both performing obstacle avoidance would learn to make their internal representations of this task similar, even though their sensors and actions differ.
(ii) Decentralized MoE Federation: Instead of aggregating entire models, agents can share task-specific MoME/MoTM/encoders/heads via peer-to-peer (P2P) communications (i.e., rather than uploading the full model to a central server, agents directly exchange only the relevant components tied to specific tasks, such as navigation or manipulation, enabling decentralized collaboration and localized knowledge transfer).
@@ IV-F Yielding Safety, Privacy, and Personalization
To make agents comply with safety, privacy, and user-specific constraints — see Fig. 3(c) — we envision the following:
(i) Module-Aware Differential Privacy: To prevent data leakage (e.g., via data inference attacks applied on the transmitted model updates from agents to the server), we envision module-aware differential privacy: a more fine-grained approach to privacy in FFMs, which applies non-uniform noise across different modules, such as encoders, heads, or MoME/MoTEs, before agents transmit them for aggregation. In this approach, the level of noise can be adjusted based on each module’s importance and usage pattern. For example, an encoder capturing generalizable features, or a frequently activated expert for sensitive tasks, would receive stronger noise, while task-specific heads or rarely used experts would require less, as they pose lower privacy risks.
(ii) Federated Behavioral Shadow Buffers: To ensure that model updates from the server are safely tested before being used by agents, we envision first deploying these updates in a “shadow mode” instead of immediately applying them, where the updated model makes predictions but its actions are not executed. The agent compares these predictions to those from its old model and looks for any big differences, while also paying attention to user reactions (e.g., corrections or feedback). This helps the agent and the system check if the new updates improve performance without causing safety issues or unwanted behavior. This shadowing process can be personalized for each user or task and activated only in sensitive situations, such as when robots take actions in shared spaces with people. Note that another option is to test the new model’s actions in a virtual environment, e.g., a digital twin, to safely observe the model’s action consequences without real-world risks.
TABLE II: Key trade-offs in FMM design for embodied AI (framed under EMBODY Dimensions). Letters ‘E’, ‘M’, ‘B’, ‘O’, ‘D’, and ‘Y’ in the last column refer to various dimensions of EMBODY.
Trade-off Dimension Option A (Advantage) Option B (Advantage) Key Tension / Real-World Implication EMBODY Dimension
Personalization vs. Global Consistency Strong per-agent adaptation via fine-tuning on-device modules (adapters, prompts, experts) Stable cross-agent coordination through shared frozen backbones Over-personalization may cause over-fitting; overly generic models may underfit unique agent dynamics E, M, O, D, Y
Communication Frequency vs. Model Quality Infrequent updates save energy and bandwidth Frequent updates improve convergence and model quality Infrequent updates can lead to low model quality; frequent updates lead to a higher resource overhead M, B, O
Privacy Assurance vs. Model Quality Strong privacy via adding a high differential privacy (DP) noise Full gradient/data sharing without adding DP noise, leading to a higher quality model Privacy constraints may suppress the model quality; Full gradient sharing risks data inference attacks M, Y
Low-overhead Modular Updates vs. Extensive Fine-tuning Light updates of the adapters, subset of MoEs, and prompts enable low-cost local tuning Extensive updates of the adapters, subset of MoEs, and prompts unlock a better learning capacity Light modular tuning aids edge deployment over resource-constrained agents but may limit the model performance E, M, B, O, Y
Safety Guarantees vs. Exploration Agility Risk-averse model updates reduce unsafe actions and enable certification Fast adaptation without risk-considerations improves exploration and model agility Safety conservatism may stall exploration; aggressive updates may introduce unsafe behavior O, D, Y
TABLE III: Benchmarking FFMs with existing embodied AI datasets: Dataset characteristics and their offered modalities and tasks [15].
Dataset / Simulator Supported Tasks Benchmark Modality Types FFM Evaluation Goals
Habitat Navigation, object search, instruction following RGB, depth, egomotion Evaluate cross-embodiment generalization, modular task head updates, zero-shot inference
Meta-World Robotic manipulation RGB, proprioception, joint Force/Torque Assess leave-one-embodiment-out generalization, task head reuse, modular fine-tuning
ManiSkill Dextrous Hand, Mobile Manipulation RGB, depth, proprioception, Egomotion Test robustness to missing/noisy modalities, cross-modal recovery via expert cascades
iGibson / RoboSuite Object Manipulation, Navigation, Household Activities RGB, depth, semantic segmentation Context-aware module scheduling, evaluation of safety-aware updates and behavioral shadowing
BEHAVIOR-1K cleaning tasks, cooking activities, organizational tasks Visual inputs, semantic segmentation, object states Benchmark continual learning and adaptation across diverse goals and environments
@@ V Evaluation Dimensions of FFMs: An EMBODY-Aligned Perspective
We next envision an evaluation framework developed around the EMBODY dimensions.
1. Embodiment Heterogeneity: To assess generalization across diverse robot morphologies and sensor configurations, models should be evaluated under “leave-one-embodiment-out” settings. This involves training on $N-1$ embodiments and deploying zero-shot on the $N^{\text{th}}$. Key metrics encompass task success rates, policy stability under various morphologies, and safety violations. Benchmark environments should simulate variable kinematics, sensor fidelity, and actuation delay.
2. Modality Richness and Imbalance: FFMs should be robust to missing or sparse modalities across agents. Evaluations should thus measure performance degradation under modality dropout, imbalance, or late-fusion timing mismatches (i.e., cases where different sensory modalities, such as vision, touch, or proprioception, arrive asynchronously or fused with misaligned timestamps due to differing sensor update rates or processing delays). To evaluate model’s robustness, the following metrics can be envisioned: modality ablation accuracy (i.e., the model’s performance when one or more input modalities are intentionally removed/masked), cross-modal transfer ability (i.e., the extent to which knowledge learned from one modality, such as vision, can be used to infer or compensate for another, such as audio or touch), and reliance bias (i.e., over-dependence on a single modality).
3. Bandwidth and Compute Constraints: FFMs deployed on resource-constrained agents must be assessed for communication and energy efficiency, under various choices of modular training/aggregation (e.g., varying numbers of involved experts, transmitting encoders vs task heads vs experts for aggregation). Evaluation should log model update size, compute per-task completion, energy-to-performance ratio, model aggregation latency, and processing depth-related metrics (e.g., percentage of activated experts during model updates and execution).
4. On-Device Continual Learning: Continual learning capability of the FFMs should be evaluated by measuring the model adaptation speed and rates of knowledge forgetting over changing environments. Metrics include forward and backward transfer (i.e., the ability to leverage prior knowledge to learn new tasks, and the ability to retain old knowledge after learning new tasks), time-to-adapt (i.e., how quickly an FFM improves performance on a new task/environment), and retention curves (i.e., performance trajectories that track how well previously learned tasks are maintained over time in the model).
5. Distributed Control and Autonomy: FFMs should be evaluated for coordination effectiveness, latency of global/local model updates, and cross-agent generalization. Metrics may include behavior divergence across agents (i.e., the degree to which different agents trained under the same FFM develop inconsistent or conflicting behaviors), emergent behavior alignment (i.e., how well the trained agents converge on coordinated or complementary behaviors during interaction with their environments), and robustness to asynchronous updates (i.e., the model’s ability to maintain stable performance when agents receive and apply model updates at different times).
6. Yielding Safety, Privacy, and Personalization: Evaluation must explicitly capture FFMs’ safety violations, such as collisions, misclassifications of dangerous zones, or regulatory compliance lapses. Also, privacy-preserving utility should be measured by how techniques such as differential privacy applied across various model modules can affect downstream embodied AI task performance. Further, personalization should be assessed in terms of task performance gain per user/agent (i.e., how much an individual agent’s task performance improves after local adaptation) and misalignment risks (i.e., the potential for personalized updates to diverge from global objectives or introduce unintended behaviors).
@@
Design Tradeoffs and Benchmarking: To operationalize FFMs in embodied AI, it is critical to understand the multi-dimensional trade-offs that emerge when balancing model personalization capability, cross-agent consistency, resource constraints, safety, and interoperability. To facilitate comprehension, in Table II we present a structured view of these trade-offs, framed through the lens of the EMBODY dimensions. Each row highlights a tension between two competing design choices, such as personalization vs. global generalization, or modular updates vs. full fine-tuning, and articulates their real-world implications. This table can serve as a guide for system designers to tailor FFM architectures and protocols according to the operational priorities and constraints of their target environments. Further, in Table III we provide a structured description of key embodied AI benchmarks that can be used to evaluate FFMs across diverse tasks and modalities.
VI Conclusion

### v2

@@ III Use Cases of FFMs in Embodied AI and EMBODY Dimensions
III-A Use Cases of FFMs in Embodied AI
We next provide three examples on the use cases of FFMs in embodied AI:
  1. 1.
    Smart Factories: FFMs can empower cooperative robots with diverse sensors, actuators, and policies to adapt to dynamic workflows and human collaborators. A typical workflow involves robotic perception and data gathering (e.g., via vision or tactile inputs), analysis and policy refinement using task‑specific modules (e.g., using a task head that outputs predictive maintenance schedules), and execution of coordinated actions in real-time. Performance requirements in such an environment include sub-second response time, low task failure rate, and high sample efficiency for adapting to new tasks (i.e., fast adaptations using few data samples).
  2. 2.
    Domestic Environments: FFMs enable assistive robots to learn and personalize to users’ routines, preferences, and spaces. A typical workflow may involve preparing morning coffee, cleaning while the user is away, restocking items when supplies run low, and reminding of evening tasks, while adapting each action to evolving household habits through on-device continual learning. Key performance targets in such settings include low computation footprint of model updates (i.e., light adaptations), continual learning without catastrophic forgetting, and high user satisfaction with minimal supervision.
  3. 3.
    Immersive XR Systems: FFMs allow distributed headsets and wearables to collaboratively fine-tune models for body-language, gaze, and gesture recognition. A typical workflow may involve capturing hand gestures, gaze, and speech during an XR session, rendering real-time feedback such as object manipulation or contextual overlays, and periodically syncing refined modality encoder modules across devices to improve recognition accuracy. Critical performance metrics include sub-second inference latency, high gesture recognition accuracy, and seamless adaptation to new users and environments within a few interactions.
$\star$ Unlike traditional FMs that often require centralized retraining, FFMs support decentralized training and adaptation in all the above scenarios, thereby preserving proprietary/user-specific data and complying with various privacy regulations. Further, different from conventional FMs, which are frozen after centralized pretraining and often lack situational responsiveness, FFMs unlock continuous, privacy-preserving adaptation to new users, tasks, and hardware configurations in all the aforementioned scenarios.
III-B Manifestation of EMBODY Dimensions
In the above scenarios, the EMBODY dimensions naturally emerge, underscoring the critical challenges that naive implementations of FFMs in embodied AI cannot address.
@@ IV Open Research Directions: An EMBODY-Aligned Agenda for FFMs
We next revisit the aforementioned EMBODY dimensions, aiming to tailor a series of open research directions (ORD). The directions are intentionally framed at a high-level to allow for diverse interpretations and encourage innovation in this underexplored domain.
Figure 3: Visualizations of the envisioned future research directions. (a): On-Device Continual Learning. Left Plot: Drift-Aware Module Creation. Right Plot: Temperature-Aware Module Aggregations. (b): Distributed Control and Autonomy. Left Plot: Federated Multi-Agent Contrastive Alignment. Right Plot: Decentralized Federation Topologies. (c): Yielding Safety, Privacy, and Personalization. Left Plot: Module-Aware Differential Privacy. Right Plot: Federated Behavioral Shadowing Buffers.
@@ IV-A Embodiment Heterogeneity
The embodied heterogeneity, caused by agents operating in various physical environments, poses a challenge to building shared FFMs that generalize across embodiments (e.g., can be trained on one or more types of robots and still perform effectively when deployed on others). To address this, we pose the following overarching ORD (see Fig. 2(a)):
ORD 1: Embodiment-Aware Information Alignment Mechanisms: We envision modules at the early layers of the model (e.g., at the MoME layers), which are augmented with locally learned tokens that encode hardware traits (e.g., kinematic range, actuator count, sensor precision). These tokens can then act as soft keys for module activations (e.g., expert selection and activation), enabling intelligent information routing decisions inside the model that control the flow of data processing and adapt it to the physical context during inference and model personalization. This approach, which we refer to as Embodiment-Conditioned Information Routing, addresses embodiment heterogeneity by enabling fine-grained, context-aware model specialization across agents with diverse morphologies without requiring hardcoded rules or full-model retraining. As a complementary approach, to avoid agents with different embodiments from diverging in representation space (i.e., variations in hardware, sensing, or control may lead to different internal encodings of similar tasks), we envision the design of consistency loss functions between related tasks performed by heterogeneous embodiments, which can be conceptualized as Cross-Embodiment Consistency Regularization, to make a closer connection between similar tasks performed across different agents.
@@ IV-B Modality Richness and Imbalance
To handle modality richness and imbalance of agents, caused by the (temporal) variations of their sensory inputs, we pose the following ORD (see Fig. 2(b)):
ORD 2: Cross-Agent Modality Compensation Strategies: We envision clustering agents with similar tasks while ensuring that the combined modalities within each cluster span a broader spectrum than any single agent can provide. Within these clusters, the server can perform federated aggregation across selected modules (e.g., MoMEs), enabling agents to benefit from peers with shared objectives as well as those that expand their sensing capabilities. This approach, which we refer to as Modality-Aware Clustered Training, allows modules within an agent’s model (e.g., MoMEs) to uncover hidden structures and relationships between modalities by leveraging knowledge from agents with richer modality sets. As an additional strategy, to address missing or degraded data streams across agents, we envision Cross-Modal Recovery via Module Cascades, where a series of specialized modules (e.g., within the MoME layer) activate only when certain modalities are unavailable or unreliable. These specialized modules should be trained in a federated manner to allow knowledge sharing across these modules from different agents, which increases their generalizability. This way, agents can collaboratively learn to fill in their missing modalities (e.g., predicting touch input from vision and movement sensors).
@@ IV-C Bandwidth and Compute Constraints
To consider the fact that many embodied agents operate on edge processors with strict energy/compute and bandwidth limitations, we pose the following ORD (see Fig. 2(c)):
ORD 3: Resource-Aware Module Usage and Update: We envision resource‑aware edge caching for FFMs, where servers (ranging from cloudlets and edge servers in mobile edge computing deployments, eNodeB base stations with local computing capabilities in cellular networks, road‑side units in vehicular edge computing scenarios, to gateway nodes in generic fog computing architectures) cache/store the modules received from agents during model aggregations, while adhering to their memory constraints. Agents can query their nearest server for modules (e.g., MoEs or adapters) relevant to their current task and environment (e.g., “following the user in a hallway”), enabling module download and reuse with only lightweight fine‑tuning. This approach, which we refer to as Context-Aware Module Caching, when accompanied by resource allocation strategies (e.g., bandwidth allocation and uplink/downlink transmit power control), can enable resource-efficient module usage in bandwidth‑ and memory‑limited wireless edge environments. As an auxiliary method, we envision Event‑Triggered Module Scheduling, where agents selectively update or train specific local modules (e.g., encoders, adapters, or MoEs) only when substantial task shifts or performance degradations occur. For instance, a mobile robot may refresh its navigation task head only when entering a new environment or encountering unfamiliar obstacles to conserve communication/computation resources.
@@ IV-D On-Device Continual Learning
To handle on-device continual learning of embodied agents, caused by facing evolving environments, user preferences, and task definitions — see Fig. 3(a) — we pose the following ORD:
ORD 4: Temporally-Tuned Module Creation and Aggregation: We envision managing task drift in FFMs, where changes in task characteristics over time may require adjusting the model architecture. At the agent level, significant task drift may trigger the fine‑tuning or splitting of a local task head into more specialized variants (e.g., a home robot’s generic “cleaning” task head evolving into distinct “kitchen cleaning” and “bedroom cleaning” heads). This approach, which we refer to as Drift‑Aware Module Creation, enables the creation of new task heads at the global model when drift patterns are common across multiple agents, while isolating the task head creation only to an agent’s local model when the drift is local. As an accompanying approach, we envision Temperature‑Aware Module Aggregation, where the server adjusts the aggregation weight and frequency of a module based on the module’s temperature, reflecting its real-time readiness or maturity. Well‑trained stable modules (i.e., “warm” modules) may be frozen and aggregated less often to reduce communication overhead, while newer or low-performing ones (i.e., “cold” modules) can be aggregated more frequently to accelerate their convergence.
TABLE II: EMBODY dimensions and their related performance metrics (Part 1); Key trade-offs in FMM design for embodied AI (Part 2); Benchmarking FFMs with existing embodied AI datasets (Part 3).
@@ IV-E Distributed Control and Autonomy
To enable embodied agents to adapt to their local tasks and contexts in a coordinated yet decentralized manner — see Fig. 3(b) — we pose the following ORD:
ORD 5: Advanced Cross-Agent Model Refinement Techniques: We introduce the concept of cross-agent refinement of FFMs through complementary strategies at both the learning and network levels. At the learning level, we envision Federated Multi‑Agent Contrastive Alignment with the ultimate goal of aligning the latent task representations of agents that perform similar tasks under different environments, hardware configurations, or user preferences. This can be achieved by using a contrastive loss objective to pull together internal representations of similar tasks and push apart those of unrelated tasks across the agents (e.g., aligning the obstacle‑avoidance task representations of an aerial drone and a ground robot). At the network level, we envision Decentralized Federation Topologies, where agents exchange their modules (e.g., MoMEs, MoTMs, adapters, encoders, or task heads) via low-power, short-range peer‑to‑peer (P2P) communication to reduce the reliance on resource-intensive uplink transmissions to the central server. Depending on agent proximity and communication topology, this can follow fully decentralized deployments, where all model exchanges occur through P2P links without any uplink usage (suitable for small‑scale, fully connected networks), or semi‑decentralized deployments, where local P2P exchanges are combined with occasional uplink transmissions from selected agents to merge knowledge across distant agents. Notably, such P2P module exchanges can be guided by mutual‑trust evaluation mechanisms between agents, a topic of long‑standing research [15].
@@ IV-F Yielding Safety, Privacy, and Personalization
To make agents comply with safety, privacy, and user-specific constraints — see Fig. 3(c) — we pose the following ORD:
ORD 6: Techniques for Module-Level Privacy and Fault Resilience: We envision strengthening privacy and operational robustness in FFMs under the presence of malicious entities in the network and/or non-ideal connectivity conditions (e.g., network jitter and packet loss). At the privacy level, we envision Module‑Aware Differential Privacy mechanisms that apply fine‑grained, non‑uniform noise to different modules (e.g., encoders, task heads, or MoME/MoTE experts) before transmission for aggregation to obscure the innate information that is encoded in the model parameters. Specifically, the noise intensity must be tuned based on each module’s characteristics; for example, modality encoders that process sensitive visual or audio streams would receive an amplified injected noise to avoid the possibility of sensitive information recovery. This can be further complemented by module‑level functional encryption (e.g., homomorphic encryption), enabling aggregation of encrypted module parameters without revealing their raw values. At the fault‑tolerance level, we propose Federated Behavioral Shadow Buffers, where model updates from the server, which might be distorted by network jitter and packet loss, are first deployed in a “shadow mode” that generates predictions without enacting actions. Agents compare these predictions to those from prior models, monitor deviations, and incorporate user feedback or corrections before applying the updates to their local models. This validation can also occur in a virtual environment (e.g., digital twin) to assess model outputs without real‑world risks.
Figure 4: Test accuracy (averaged over the two tasks) versus cumulative energy consumption across users (left subplot) and overall training latency (right subplot) for different methods. The “Convergence Point ($\pm 0.5\%$)” markers denote the energy level and latency at which each method converges and oscillates with only less than $\pm 0.5\%$ in accuracy. FFM + P2P Relay consistently delivers faster convergence while converging to the same final accuracy as FFM, underscoring the role of P2P communication among embodied AI agents in minimizing energy usage and latency. Isolated local training, where embodied AI agents do not engage in cross-embodiment/user knowledge sharing, converges to significantly lower accuracy.
@@ V Evaluation Dimensions of FFMs: An EMBODY-Aligned Perspective
We next envision an evaluation framework developed around the EMBODY dimensions, structured into three key components as outlined in the overarching Table II. In particular, in Part 1 of this table, we present the six EMBODY dimensions and their brief descriptions, followed by the corresponding performance metrics that can be used for their evaluation. For each dimension, we also provide an interpretation of the reported metrics, explaining what they reveal about system behavior and robustness. Further, we discuss the real-world implications of addressing each EMBODY dimension. Next, we consider the fact that to operationalize FFMs in embodied AI, it is critical to understand the multi-dimensional trade-offs that are rooted in the evaluation of EMBODY dimensions. To facilitate comprehension, in Part 2 of Table II, we present a structured view of these trade-offs. Each row highlights a tension between two competing design choices, such as personalization vs. global generalization, and articulates their real-world implications. This table can serve as a guide for system designers to tailor FFM architectures and protocols according to the operational priorities and constraints of their target environments. Finally, in Part 3 of Table II, we provide a structured description of key embodied AI datasets that can be used to evaluate FFMs across diverse tasks and modalities.
VI Prototype of FFMs
We consider an edge network comprising $35$ embodied AI robots. The edge robots are partitioned into $7$ clusters, where clusters represent different zones/areas ($5$ robots are considered in each cluster, which can further form P2P networks as in ORD 5). We presume a scenario where the embodied AI robots aim to learn two heterogeneous tasks/datasets: (i) learning to engage in question-answering (e.g., for answering queries about everyday environments, such as identifying what objects are present, what actions are happening, or where items are located) via recognizing generic images and the description of objects in them via GQA dataset (https://cs.stanford.edu/people/dorarad/gqa/download.html), and (ii) recognizing different shopping products and their descriptions (e.g., for assisting individuals/customers, particularly those with visual impairments, with shopping) through VizWiz dataset (http://vizwiz.org/tasks-and-datasets/vqa/), using two data modalities (text and image).
Our implementations are publicly available on GitHub: https://github.com/payamsiabd/M3T-FFM-EmbodiedAI, detailing the P2P network topology and uplink/downlink channel models. The non-iid distribution of each dataset/task across agents follows Dirichlet distribution [5] with concentration parameter $0.5$. We adopt ViLT (with the size of 328 MB) as the backbone of the FMs deployed/trained on robots, which offers two advantages: (i) it employs a lightweight text embedding layer instead of the LLM–based text encoder used in VauLT (Fig. 1(d)), and (ii) it is designed for image–text multimodality, which matches the modalities present in our datasets.
We fine‑tune the model using lightweight adapters embedded within every transformer layer alongside task‑specific output heads (with the total size of 6MB). Each adapter adopts a bottleneck architecture with a hidden dimension of $256$, comprising a down‑projection layer, a GELU nonlinearity, and an up‑projection layer.
We examine the following methods: (i) Local FM Deployment: Each robot trains its own FM without aggregating/knowledge sharing with other robots. (ii) FFM Deployment: After one epoch of local training, each robot sends its fine-tuned adapter and task head module parameters to the cloud, which conducts a module aggregation using FedAvg [5]. (iii) FFM + P2P Module Relaying: In each cluster, a randomly selected robot serves as the cluster head and performs the following operations. After completing local training, all robots in the cluster transmit their fine-tuned parameters to the cluster head via multi‑hop P2P links along the shortest path, where the parameters are aggregated. The cluster head then solely uploads its aggregated parameters to the cloud for global aggregation, thereby reducing dependence on resource‑intensive uplink transmissions.
In Fig. 4, we capture various performance metrics, such as energy-to-performance ratio (left subplot) and latency-to-performance ratio (right subplot), where the performance is measured via the test accuracy. Comparing the performance of the ‘Local’ baseline with other FFM variants highlights the role of global knowledge sharing in FFMs, where the local model’s performance saturates to a low accuracy. Further, observing the plots reveals the energy usage and latency of FFM fine-tuning processes. Specifically, the FFM + P2P Relay achieves the same converged accuracy as its FFM counterpart, attributable to replicating model aggregations via multi‑hop P2P model exchanges, while incurring lower energy consumption and latency, mainly due to the efficient communications enabled by low‑cost, short‑range P2P links.
VII Conclusion